OpenStudy (anonymous):

p+9/5=q for p show work please

OpenStudy (saifoo.khan):

\[\frac{p+9}{5}=q\]\[p+9 = 5q\]

OpenStudy (saifoo.khan):

p = 5q - 9

OpenStudy (anonymous):

is tht the answer ?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

ohkay thx

OpenStudy (saifoo.khan):

np